OpenStudy (anonymous):

How to find the solution set of √(x+4)+√(9-x)=5

OpenStudy (anonymous):

Start by squaring everything

OpenStudy (anonymous):

how was it? :))

OpenStudy (anonymous):

What do you mean? Square EVERYTHING: So do: \[(\sqrt{x+4})^2 + (\sqrt{9-x})^2 = 5^2\]

OpenStudy (anonymous):

Then just solve for x.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem. mind marking best response?

OpenStudy (anonymous):

How will I solve this "A 30% of barium chloride is mixed with 10 liters of water to form 20% solution. How many Liters of original solution did we start with?"